Which of the following fractions have terminating decimals and which have repeating decimals.
The fractions are:
1/2, 1/3, 1/4, 1/5, 1/6, 1/7, 1/8, 1/9, 1/10, 1/11, 1/12
Show Decimal Form
" Which are Terminate
and Prime Factorization of the Denominator
I am very lost on this.
Thanks if you can help
Sher

Answer by checkley71(8403)

TO FIND THE DECIMAL EQUIVALENT OF A FRACTION JUST DIVIDE THE NUMERATOR BY THE DENOMINATOR THUS
1/2=.5 1/3=.33 1/3, 1/4=.25,1/5=.20,1/6=.16 2/3,1/7=.14 2/7,1/8=.125,
1/9=.11 1/9, 1/10=.1,1/11=.09 1/11,1/12=.08 1/3